# Gromov-Wasserstein Barycenters: The Analysis Problem

### v1

@@ 5.2. Classification
Figure 8. t-SNE of estimated $\lambda$ coordinates. One-shot clustering accuracy (K-Means): 0.9, achieved using a single template per class for both (A) Algorithm 2 and (B) Algorithm 4.
a Three templates per class.
b Six templates per class.
Figure 9. t-SNE of estimated $\lambda$ coordinates via Algorithm 2. Multiple templates per class: (A) 3, (B) 6. Clustering accuracy (K-Means): (A) 0.95, (B) 0.98.
@@ 6. Conclusions
This paper addresses the problem of recovering barycentric coordinates in the context of finite metric spaces, or more generally, networks with weighted dissimilarities between nodes, embedded via the Gromov–Wasserstein (GW) distance, referred to as the GW analysis problem. While the literature offers at least two established methods for computing GW barycenters, this work builds upon them and proposes two corresponding solutions to the GW analysis problem, each with its own algorithm. Moreover, we show that these two perspectives are not entirely independent: under suitable parameter choices, they share common ground. Notably, both proposed solutions correspond to critical points of convex functions (Theorems 3.6 and 4.7). Finally, we introduce strategies for estimating barycentric coordinates that may remain useful even when the target object is not an exact GW barycenter.
This paper focuses on the case of discrete metric measure spaces (or discrete networks), although the GW framework also applies in the continuum setting. Extending our fixed-point and gradient-based approaches to a broader context is of interest. In addition, this paper does not establish a full equivalence between Karcher means and GW barycenters. In the case of classical OT, such an equivalence can be shown under certain assumptions [68]. Developing analogous conditions in the GW setting remains an open direction for future research.
@@ Appendix A Background: The Barycentric Coding Model (BCM)
$\mathbf{x}=\sum_{s\in[S]}\lambda_{s}\mathbf{x}^{s},\qquad\text{ for some }\lambda=(\lambda_{1},\dots,\lambda_{S})\in\Delta_{S-1}.$ (41)
@@
$\Delta_{S-1}\ni\lambda\longmapsto\mathbf{x}_{\lambda}:=\sum_{s\in[S]}\lambda_{s}\mathbf{x}^{s}\in\mathrm{B}.$ (42)
@@
$\Delta_{S-1}\ni\lambda\longmapsto\mathbf{x}_{\lambda}=\operatorname*{arg\,min}_{\mathbf{x}\in\mathbb{R}^{n}}\sum_{s\in[S]}{\lambda_{s}}\|\mathbf{x}^{s}-\mathbf{x}\|^{2}\in\mathrm{B}.$ (43)
@@
$\mathbf{x}\longmapsto\lambda_{\mathbf{x}}:=\operatorname*{arg\,min}_{\lambda\in\Delta_{S-1}}\|\mathbf{x}-\sum_{s\in[S]}\lambda_{s}\mathbf{x}^{s}\|^{2}.$ (44)
Moreover, one can extend (44) by considering $\mathbf{x}\in\mathbb{R}^{n}$ instead of restricting to $\mathbf{x}\in\mathrm{B}$. This amounts to projecting $\mathbf{x}$ onto $\mathrm{B}$ and retaining only the coordinates $\lambda_{\mathbf{x}}$ of the projected point181818 Let $\Phi:\mathbb{R}^{S}\to\mathbb{R}^{n}$ be the matrix whose columns are the vectors $\mathbf{x}^{1},\dots,\mathbf{x}^{S}$, and consider the quadratic function $f(\lambda):=\|\mathbf{x}-\sum_{s\in[S]}\lambda_{s}\mathbf{x}^{s}\|^{2}=\|\mathbf{x}-\Phi\lambda\|^{2}$ for all $\lambda\in\Delta_{S-1}$. Since $f$ is continuous and the simplex is compact, a minimizer $\lambda_{\mathbf{x}}$ for problem (44) exists. Define $\mathbf{z}:=\Phi\lambda_{\mathbf{x}}\in\mathrm{B}$. By construction and since $\mathrm{B}=\Phi(\Delta_{S-1})$, we have $\|\mathbf{x}-\mathbf{z}\|=\min_{\lambda\in\Delta_{S-1}}\|\mathbf{x}-\Phi\lambda\|=\min_{\mathbf{y}\in\mathrm{B}}\|\mathbf{x}-\mathbf{y}\|$. Thus, by convex analysis, this last property characterizes $\mathbf{z}$ as the unique orthogonal projection of $\mathbf{x}$ onto $\mathrm{B}$ (even if $\lambda_{\mathbf{x}}$ is not unique). ; consequently, the orthogonal projection of $\mathbf{x}\in\mathbb{R}^{n}$ onto the closed convex set $\mathrm{B}$ is given by $\sum_{s\in[S]}(\lambda_{\mathbf{x}})_{s}\,\mathbf{x}^{s}$ [60]. Moreover, problem (44) can be rewritten as
$\mathbf{x}\longmapsto\lambda_{\mathbf{x}}:=\operatorname*{arg\,min}_{\lambda\in\Delta_{S-1}}\|\mathbf{x}-\mathbf{x}_{\lambda}\|^{2},$ (45)
@@
$\Delta_{S-1}\ni\lambda\longmapsto\mathbf{x}_{\lambda}\in\operatorname*{arg\,min}_{\mathbf{x}\in\mathrm{G}}\sum_{s\in[S]}\lambda_{s}\,dist^{2}(\mathbf{x}^{s},\mathbf{x}).$ (46)
@@
$\mathrm{G}\ni\mathbf{x}\longmapsto\lambda_{\mathbf{x}}\in\operatorname*{arg\,min}_{\lambda\in\Delta_{S-1}}d^{2}(\mathbf{x},\mathbf{x}_{\lambda}),$ (47)
@@ B.2. GW Synthesis Problem – Fixed-Point Iteration Scheme
In this section we consider the GW synthesis problem (9). First, for completeness, we provide a proof of Proposition 3.2 (see also [46, Prop. 3]), followed by several remarks.
@@
$0\leq J_{\lambda}(\mathbf{Y})$ $=\sum_{s\in[S]}\lambda_{s}\sum_{i\in[N^{s}],j\in[M]}\sum_{k\in[N^{s}],l\in[M]}|\mathbf{X}_{ik}^{s}-\mathbf{Y}_{jl}|^{2}\pi_{ij}^{s}\pi_{kl}^{s}$ (48)
@@
$\qquad-2\sum_{s\in[S]}\sum_{i\in[N^{s}],j\in[M]}\sum_{k\in[N^{s}],l\in[M]}\lambda_{s}\mathbf{X}_{ik}^{s}\mathbf{Y}_{jl}\pi_{ij}^{s}\pi_{kl}^{s}.$ (49)
@@
$\frac{\partial\,J_{\lambda}}{\partial\mathbf{Y}_{j_{0}\,l_{0}}}=2\,\mathrm{q}_{j_{0}}\mathrm{q}_{l_{0}}\mathbf{Y}_{j_{0}\,l_{0}}-2\sum_{s\in[S]}\lambda_{s}\sum_{i,k\in[N^{s}]}\mathbf{X}_{ik}^{s}\pi_{ij_{0}}^{s}\pi_{kl_{0}}^{s}.$ (50)
@@
In general, since $\mathrm{q}_{j}\geq 0$ for all $j$, the Hessian is positive semidefinite, and therefore, by convexity and the fact that the objective function is nonnegative, a minimizer exists. Moreover, if $\mathrm{q}_{j}>0$ for all $j$, the Hessian is positive definite, implying that the problem admits a unique global minimum. The entries of the minimum $\mathbf{Y}^{*}\in\mathbb{R}^{M\times M}$ are characterized by setting the gradient expression (50) to zero, yielding the closed-form solution (13).
@@
Notice that if $\{\mathbf{X}^{s}\}_{s\in[S]}$ are symmetric, then formula (13) defines a symmetric matrix $\mathbf{Y}^{*}$. Similarly, if the matrices $\{\mathbf{X}^{s}\}_{s\in[S]}$ have non-negative entries or are positive semidefinite, then $\mathbf{Y}^{*}$ will also have non-negative entries or be positive semidefinite [46, Prop. 3]. However, in general, $\mathbf{Y}^{*}$ does not have zero-trace (or zero-diagonal), even if the template matrices $\{\mathbf{X}^{s}\}_{s\in[S]}$ have zero trace. Therefore, even if the templates are distance matrices, $\mathbf{Y}^{*}$ given by (13) does not necessarily define a distance matrix.
@@
$\geq k_{1}\|\mathbf{Y}\|_{\mathrm{Frob}}^{2}-k_{2}\|\mathbf{Y}\|_{\mathrm{Frob}}$ (51)
@@
$\mathbf{X}=\frac{1}{\mathrm{p}\mathrm{p}^{T}}\odot(\pi^{*})^{T}\mathbf{X}\pi^{*}\qquad\forall\pi^{*}\in\Pi(\mathrm{p},\mathrm{p})\text{ optimal for }GW((\mathbf{X},\mathrm{p}),(\mathbf{X},\mathrm{p})).$ (52)
[…]
@@ B.3. Weak Isomorphisms
Let $\mathbb{X},\mathbb{Y}$ be two mm-spaces. By [38, Thm. 5.1] $GW(\mathbb{X},\mathbb{Y})=0$ if and only if $\mathbb{X}$ and $\mathbb{Y}$ are strongly isomorphic. In more generality, if $\mathbb{X},\mathbb{Y}$ are two networks, then $GW(\mathbb{X},\mathbb{Y})=0$ if and only if $\mathbb{X}$ and $\mathbb{Y}$ are weakly isomorphic [12, Thms. 2.3 and 2.4]. As a consequence, two mm-spaces are strongly isomorphic if and only if they are weakly isomorphic. To gain some intuition we provide a proof for the finite case. The proof of the general case follows the same ideas but requires some technical results such as [38, Lem. 2.2 or Lem. 10.4].
@@
$d_{X}(\varphi_{X}(z),\varphi_{X}(z^{\prime}))=d_{Y}(\varphi_{Y}(z),\varphi_{Y}(z^{\prime}))\qquad\forall z,z^{\prime}\in Z.$ (53)
@@
As an informal interpretation, under weak isomorphism two nodes $y,y^{\prime}\in Y$ are ‘the same’ if they have the same internal connectivity, i.e., $\omega_{Y}(y,y)=\omega_{Y}(y,y^{\prime})=\omega_{X}(y^{\prime},y)=\omega_{Y}(y^{\prime},y^{\prime})$, and the same external connectivity, i.e., $\omega_{Y}(y,a)=\omega_{Y}(y^{\prime},a)$ and $\omega_{Y}(a,y)=\omega_{Y}(a,y^{\prime})$ for all $a\in Y$ [13]. In our example, the two nodes in the space $Y$ with mass $0.3$ are ‘the same’.
@@
By using the blow-up technique (i.e., by using Algorithm 3 with an optimal coupling for $GW(\mathbb{X},\mathbb{Y})$), the logarithmic map $Log_{[\mathbb{Y}]}:\mathcal{GM}\longrightarrow\mathrm{Tan}_{[\mathbb{Y}]}$, that is, the local inverse of the exponential map (27), can be written as: $Log_{[\mathbb{Y}]}([\mathbb{X}]):=[\mathbf{X}_{b}-\mathbf{Y}_{b}]$. We refer the reader to [14, Sec. 3.4].
@@
is a GW geodesic between $[\mathbb{X}]$ and $[\mathbb{Y}]$, in the sense that $GW(\gamma(t),\gamma(s))=|t-s|GW(\mathbb{X},\mathbb{Y})$ for all $0\leq s,t\leq 1$. In particular, ${\omega_{X_{b}}}-{\omega_{Y_{b}}}$ (represented by the matrix $\mathbf{X}_{b}-\mathbf{Y}_{b}$) can be viewed as its velocity. We refer the reader to [56, 14]. As a consequence, the GW barycenter between two references $[\mathbb{X}]$, $[\mathbb{Y}]$ with uniform weights $\lambda=(1/2,1/2)$ is given by the network class $\gamma(1/2)=[(\frac{1}{2}\mathbf{X}_{b}+\frac{1}{2}\mathbf{Y}_{b},\mathrm{q}_{b})]$. Indeed, by using the triangle inequality, for every network class $[\mathbb{Z}]$ we have:
@@
In general, $\gamma(t)$ is a GW barycenter between $[\mathbb{X}]$ and $[\mathbb{Y}]$ with weights $\lambda=(1-t,t)$ because for every network class $[\mathbb{Z}]$ we have:
@@
Let us show that a minimizer $[\mathbb{Y}^{*}]$ of a differentiable functional in GW space, $G:\mathcal{GW}\to\mathbb{R}$, must satisfy the first-order optimality condition, that is, $\nabla G([\mathbb{Y}^{*}])=0$. First, by assumption we have that $G([\mathbb{Y}^{*}])\leq G([\mathbb{Y}])$ for all $[\mathbb{Y}]\in\mathcal{GW}$. In particular, $G([\mathbb{Y}^{*}])\leq G(Exp_{[\mathbb{Y}^{*}]}(tv))$ for all $[v]\in\mathrm{Tan}_{[\mathbb{Y}^{*}]}$. Thus,
@@
In particular, for the direction $[v]=-\nabla G([\mathbb{Y}^{*}])$ (which exists because tangent vectors are modeled as elements of $L^{2}$-spaces) we have
@@
Thus, the authors in [14] develop a first-order calculus on GW space–comprising directional derivatives of functionals, gradients as linear functionals (or vectors), first-order optimality conditions for minimizers, and linear approximations of functions around a point–by equipping it with a formal tangent space at each point $[\mathbb{Y}]$, based on a Riemannian-like framework. This ends the remark.
For completeness, we revisit the proof of Proposition 4.3, originally presented in [14, Prop. 8]. Before doing so, we state the following lemma and a subsequent observation.
@@
    For any $v\in L^{2}(Y\times Y,\mu_{Y}\otimes\mu_{Y})$, $v\simeq\varphi^{*}v\in L^{2}(Z\times Z,\mu_{Z}\otimes\mu_{Z})$ (i.e., $[v]=[\varphi^{*}v]\in\mathrm{Tan}_{[\mathbb{Y}]}$).

### v2

@@ 5.2. Clustering, Classification and Visualization
Figure 10. K-means clustering partitions: Comparison of decision regions on the simplex for three MNIST classes (digits 0, 4, and 7), using 200 sample point clouds per class. Algorithm 2 achieves Adjusted Rand Index $\mathrm{ARI}\approx 0.6$ and requires $\approx 405$ seconds to compute all GW-barycentric coordinates. Algorithm 4 yields $\mathrm{ARI}\approx 0.55$, requires $\approx 1293$ seconds to compute all coordinates, and has an average blow-up size ($M_{b}$): 661.8. Clear clustering for digit class 0. In contrast, digit classes 4 and 7 exhibit overlap, suggesting the presence of shared geometric shapes between them.
@@ 6. Conclusions
This paper addresses the problem of recovering barycentric coordinates in the context of finite metric spaces, or more generally, networks with weighted dissimilarities between nodes, embedded via the Gromov-Wasserstein (GW) distance, referred to as the GW analysis problem. While the literature offers at least two established methods for computing GW barycenters, this work builds upon them and proposes two corresponding solutions to the GW analysis problem, each with its own algorithm. Moreover, we show that these two perspectives are not entirely independent: under suitable parameter choices, they share common ground. Notably, both proposed solutions correspond to critical points of convex functions (Theorems 3.6 and 4.7). Our results pertain to recovery of optimal weights when a network lies exactly in the GW barycentric span. Nevertheless, our strategies are applicable, and can still be useful, when target objects are not exact barycenters. Obtaining theoretical guarantees to learn optimal weights for networks outside the span is an interesting problem. In addition to Section 5.2, where experiments are conducted on non-synthetic data, we present preliminary empirical work and a discussion of the challenges of this problem in Suppl. Mat. E.
This paper focuses on the case of discrete metric measure spaces (or discrete networks), although the GW framework also applies in the continuum setting. Extensions of our fixed-point and gradient-based approaches to a broader context could be of interest. This paper does not establish a full equivalence between Karcher means and GW barycenters. In the case of classical OT, such an equivalence can be shown under certain assumptions [72]. Developing analogous conditions in the GW setting remains an open problem for future research.
There are also several experimental directions to explore. For instance, efficient implementations of the blow-up technique, so that, when applied to a collection of networks, they could identify a smallest equivalent representative of each network while preserving global alignment (enabling compressed storage, and faster computations, by avoiding repeated values). Also, further experiments in settings with alternative node dissimilarities (e.g., replacing the Euclidean distance with intrinsic notions such as geodesic distances on manifolds) would also help solidify the applicability of the GW-BCM to real world problems. Initial experimental results with $p$-Fermat [35] and graph distances in the GW-BCM framework are in Suppl. Mat. D.3 and D.1.2.
@@ Appendix A Background: The Barycentric Coding Model (BCM)
(41) $\mathbf{x}=\sum_{s\in[S]}\lambda_{s}\mathbf{x}^{s},\qquad\text{ for some }\lambda=(\lambda_{1},\dots,\lambda_{S})\in\Delta_{S-1}.$
@@
(42) $\Delta_{S-1}\ni\lambda\longmapsto\mathbf{x}_{\lambda}:=\sum_{s\in[S]}\lambda_{s}\mathbf{x}^{s}\in\mathrm{B}.$
@@
(43) $\Delta_{S-1}\ni\lambda\longmapsto\mathbf{x}_{\lambda}=\operatorname*{arg\,min}_{\mathbf{x}\in\mathbb{R}^{n}}\sum_{s\in[S]}{\lambda_{s}}\|\mathbf{x}^{s}-\mathbf{x}\|^{2}\in\mathrm{B}.$
@@
(44) $\mathbf{x}\longmapsto\lambda_{\mathbf{x}}:=\operatorname*{arg\,min}_{\lambda\in\Delta_{S-1}}\|\mathbf{x}-\sum_{s\in[S]}\lambda_{s}\mathbf{x}^{s}\|^{2}.$
Moreover, one can extend (44) by considering $\mathbf{x}\in\mathbb{R}^{n}$ instead of restricting to $\mathbf{x}\in\mathrm{B}$. This amounts to projecting $\mathbf{x}$ onto $\mathrm{B}$ and retaining only the coordinates $\lambda_{\mathbf{x}}$ of the projected point [^11] Let $\Phi:\mathbb{R}^{S}\to\mathbb{R}^{n}$ be the matrix whose columns are the vectors $\mathbf{x}^{1},\dots,\mathbf{x}^{S}$, and consider the quadratic function $f(\lambda):=\|\mathbf{x}-\sum_{s\in[S]}\lambda_{s}\mathbf{x}^{s}\|^{2}=\|\mathbf{x}-\Phi\lambda\|^{2}$ for all $\lambda\in\Delta_{S-1}$. Since $f$ is continuous and the simplex is compact, a minimizer $\lambda_{\mathbf{x}}$ for problem (44) exists. Define $\mathbf{z}:=\Phi\lambda_{\mathbf{x}}\in\mathrm{B}$. By construction and since $\mathrm{B}=\Phi(\Delta_{S-1})$, we have $\|\mathbf{x}-\mathbf{z}\|=\min_{\lambda\in\Delta_{S-1}}\|\mathbf{x}-\Phi\lambda\|=\min_{\mathbf{y}\in\mathrm{B}}\|\mathbf{x}-\mathbf{y}\|$. Thus, by convex analysis, this last property characterizes $\mathbf{z}$ as the unique orthogonal projection of $\mathbf{x}$ onto $\mathrm{B}$ (even if $\lambda_{\mathbf{x}}$ is not unique). ; consequently, the orthogonal projection of $\mathbf{x}\in\mathbb{R}^{n}$ onto the closed convex set $\mathrm{B}$ is given by $\sum_{s\in[S]}(\lambda_{\mathbf{x}})_{s}\,\mathbf{x}^{s}$ [63]. Moreover, problem (44) can be rewritten as
(45) $\mathbf{x}\longmapsto\lambda_{\mathbf{x}}:=\operatorname*{arg\,min}_{\lambda\in\Delta_{S-1}}\|\mathbf{x}-\mathbf{x}_{\lambda}\|^{2},$
@@
(46) $\Delta_{S-1}\ni\lambda\longmapsto\mathbf{x}_{\lambda}\in\operatorname*{arg\,min}_{\mathbf{x}\in\mathrm{G}}\sum_{s\in[S]}\lambda_{s}\,dist^{2}(\mathbf{x}^{s},\mathbf{x}).$
@@
(47) $\mathrm{G}\ni\mathbf{x}\longmapsto\lambda_{\mathbf{x}}\in\operatorname*{arg\,min}_{\lambda\in\Delta_{S-1}}d^{2}(\mathbf{x},\mathbf{x}_{\lambda}),$
@@
A.1. BCM: Classical OT vs. Euclidean framework
To illustrate the qualitative differences between geometric models, we consider the problem of interpolating between two weighted point clouds representing handwritten digits. This corresponds to visualizing the barycentric space generated by two template point clouds. Using the Point Cloud MNIST 2D dataset [24], we examine how the choice of geometry influences the resulting intermediate shapes. Fig. 11 compares the interpolations obtained using Wasserstein geometry versus using standard Euclidean geometry. The reader can also compare Figure 11 with Figure 1 in the main text.
Figure 11. Illustration of different barycenter spaces between two point clouds selected from the Point Cloud MNIST 2D dataset [24], shown for various $t\in[0,1]$ with interpolation coordinates (1t,t)(1-t,t) on the horizontal axis. Top: OT interpolations (with respect to the 2-Wasserstein geometry). Bottom: Linear interpolations (with respect to the Euclidean geometry).
Analytically, consider the space of discrete probability measures with finite support on $\mathbb{R}$, equipped with the 2-Wasserstein metric $W_{2}(\cdot,\cdot)$. Let the template measures be $\delta_{0}$ and $\delta_{1}$, where $\delta_{x}$ denotes the Dirac measure centered at $x$. The corresponding barycenter space consists of the set $\{\delta_{t}:\,t\in[0,1]\}$, which is obtained by solving the synthesis problem (1) (in the main text) in this specific setting. In other words, each measure $\mu=\delta_{t}$ minimizes the functional $\mu\mapsto t\,W_{2}^{2}(\delta_{0},\mu)+(1-t)\,W_{2}^{2}(\delta_{1},\mu)$. Hence, we can interpret the coordinates of the measure $\delta_{t}$ with respect to the templates $\delta_{0}$ and $\delta_{1}$ as the vector (t,1t)(t,1-t) in $\Delta_{1}$. In this simple example, barycenters correspond to Dirac masses supported on convex combinations of the points 0 and 11–that is, Euclidean barycenters. In more general settings, however, measures or distributions have the ability to represent more complex data than individual points in space. For instance, interpolating between two templates can be understood as constructing the barycentric space generated by those reference templates. Specifically, if $\mu_{0}$ and $\mu_{1}$ are the probability distributions corresponding to two templates, and $\pi$ is an OT plan between them in the 2-Wasserstein sense, then the interpolated measure $\mu_{t}$ defined as the pushforward of $\pi$ under the map $(x,y)\mapsto(1-t)x+ty$, that is, $\mu_{t}:=((1-t)x+ty)_{\#}\pi$ for $t\in[0,1]$, coincides with the Wasserstein barycenter of $\mu_{0}$ and $\mu_{1}$ with respective weights (1t)(1-t) and $t$: $\mu_{t}=\operatorname*{arg\,min}_{\mu}\left\{(1-t)W_{2}^{2}(\mu,\mu_{0})+tW_{2}^{2}(\mu,\mu_{1})\right\}$.
Finally, we notice that one of the main differences between the 2-Wasserstein distance and the GW distance considered in this work, lies in how they support barycenter constructions. In the $W_{2}$ setting, barycenters between two or more measures can be constructed via displacement interpolation using OT maps under regularity assumptions [38, 1]. This interpolation aligns mass in a common ambient space and is underpinned by the pseudo-Riemannian structure of Wasserstein space [77]. In contrast, GW distance compares measures on different metric spaces by optimizing over structural correspondences rather than spatial displacements. While geodesics between two metric measure spaces exist in GW space and can yield meaningful barycenters (see Figure 1 in the main text), there is no analogous displacement interpolation for more than two templates. This reflects the absence of a pseudo-Riemannian framework (in the classical sense [45, 4]), limiting the direct geometric interpretation of GW barycenters.
@@ B.2. GW Synthesis Problem – Fixed-Point Iteration Scheme
In this section we consider the GW synthesis problem (6). First, for completeness, we provide a proof of Proposition 3.2 (see also [49, Prop. 3]), followed by several remarks.
@@
(48) $0\leq J_{\lambda}(\mathbf{Y})$ $=\sum_{s\in[S]}\lambda_{s}\sum_{i\in[N^{s}],j\in[M]}\sum_{k\in[N^{s}],l\in[M]}|\mathbf{X}_{ik}^{s}-\mathbf{Y}_{jl}|^{2}\pi_{ij}^{s}\pi_{kl}^{s}$
@@
(49) $\qquad-2\sum_{s\in[S]}\sum_{i\in[N^{s}],j\in[M]}\sum_{k\in[N^{s}],l\in[M]}\lambda_{s}\mathbf{X}_{ik}^{s}\mathbf{Y}_{jl}\pi_{ij}^{s}\pi_{kl}^{s}.$
@@
(50) $\frac{\partial\,J_{\lambda}}{\partial\mathbf{Y}_{j_{0}\,l_{0}}}=2\,\mathrm{q}_{j_{0}}\mathrm{q}_{l_{0}}\mathbf{Y}_{j_{0}\,l_{0}}-2\sum_{s\in[S]}\lambda_{s}\sum_{i,k\in[N^{s}]}\mathbf{X}_{ik}^{s}\pi_{ij_{0}}^{s}\pi_{kl_{0}}^{s}.$
@@
In general, since $\mathrm{q}_{j}\geq 0$ for all $j$, the Hessian is positive semidefinite, and therefore, by convexity and the fact that the objective function is nonnegative, a minimizer exists. Moreover, if $\mathrm{q}_{j}>0$ for all $j$, the Hessian is positive definite, implying that the problem admits a unique global minimum. The entries of the minimum $\mathbf{Y}^{*}\in\mathbb{R}^{M\times M}$ are characterized by setting the gradient expression (50) to zero, yielding the closed-form solution (10). ∎
@@
Notice that if $\{\mathbf{X}^{s}\}_{s\in[S]}$ are symmetric, then formula (10) defines a symmetric matrix $\mathbf{Y}^{*}$. Similarly, if the matrices $\{\mathbf{X}^{s}\}_{s\in[S]}$ have non-negative entries or are positive semidefinite, then $\mathbf{Y}^{*}$ will also have non-negative entries or be positive semidefinite [49, Prop. 3]. However, in general, $\mathbf{Y}^{*}$ does not have zero-trace (or zero-diagonal), even if the template matrices $\{\mathbf{X}^{s}\}_{s\in[S]}$ have zero trace. Therefore, even if the templates are distance matrices, $\mathbf{Y}^{*}$ given by (10) does not necessarily define a distance matrix.
@@
(51) $\geq k_{1}\|\mathbf{Y}\|_{\mathrm{Frob}}^{2}-k_{2}\|\mathbf{Y}\|_{\mathrm{Frob}}$
@@
(52) $\mathbf{X}=\frac{1}{\mathrm{p}\mathrm{p}^{T}}\odot(\pi^{*})^{T}\mathbf{X}\pi^{*}\qquad\forall\pi^{*}\in\Pi(\mathrm{p},\mathrm{p})\text{ optimal for }GW((\mathbf{X},\mathrm{p}),(\mathbf{X},\mathrm{p})).$
To reinforce (52), let us verify it by exhaustion in the case $N=M=2$, $\mathbf{X}$ anti-diagonal matrix and $\mathrm{p}$ uniform. That is, consider $\mathbf{X}=\begin{pmatrix}0&c\\ c&0\end{pmatrix}$ for some $c>0$ and $\mathrm{p}=(.5,.5)$. Then
@@ B.3. Weak Isomorphisms
Let $\mathbb{X},\mathbb{Y}$ be two mm-spaces. By [42, Thm. 5.1] $GW(\mathbb{X},\mathbb{Y})=0$ if and only if $\mathbb{X}$ and $\mathbb{Y}$ are strongly isomorphic. In more generality, if $\mathbb{X},\mathbb{Y}$ are two networks, then $GW(\mathbb{X},\mathbb{Y})=0$ if and only if $\mathbb{X}$ and $\mathbb{Y}$ are weakly isomorphic [13, Thms. 2.3 and 2.4]. As a consequence, two mm-spaces are strongly isomorphic if and only if they are weakly isomorphic. To gain some intuition we provide a proof for the finite case. The proof of the general case follows the same ideas but requires some technical results such as [42, Lem. 2.2 or Lem. 10.4].
@@
(53) $d_{X}(\varphi_{X}(z),\varphi_{X}(z^{\prime}))=d_{Y}(\varphi_{Y}(z),\varphi_{Y}(z^{\prime}))\qquad\forall z,z^{\prime}\in Z.$
@@
As an informal interpretation, under weak isomorphism two nodes $y,y^{\prime}\in Y$ are ‘the same’ if they have the same ‘internal connectivity’, i.e., $\omega_{Y}(y,y)=\omega_{Y}(y,y^{\prime})=\omega_{X}(y^{\prime},y)=\omega_{Y}(y^{\prime},y^{\prime})$, and the same ‘external connectivity’, i.e., $\omega_{Y}(y,a)=\omega_{Y}(y^{\prime},a)$ and $\omega_{Y}(a,y)=\omega_{Y}(a,y^{\prime})$ for all $a\in Y$ [14]. In our example, the two nodes in the space $Y$ with mass $0.3$ are ‘the same’.
@@
By using the blow-up technique (i.e., by using Algorithm 3 with an optimal coupling for $GW(\mathbb{X},\mathbb{Y})$), the logarithmic map $Log_{[\mathbb{Y}]}:\mathcal{GM}\longrightarrow\mathrm{Tan}_{[\mathbb{Y}]}$, that is, the local inverse of the exponential map (26), can be written as: $Log_{[\mathbb{Y}]}([\mathbb{X}]):=[\mathbf{X}_{b}-\mathbf{Y}_{b}]$. We refer the reader to [15, Sec. 3.4].
@@
is a GW geodesic between $\mathbb{X}]$ and $\mathbb{Y}]$, in the sense that $GW(\gamma(t),\gamma(s))=|t-s|GW(\mathbb{X},\mathbb{Y})$ for all $0\leq s,t\leq 1$. In particular, ${\omega_{X_{b}}}-{\omega_{Y_{b}}}$ (represented by the matrix $\mathbf{X}_{b}-\mathbf{Y}_{b}$) can be viewed as its velocity. We refer the reader to [59, 15]. As a consequence, the GW barycenter between two references $\mathbb{X}]$, $\mathbb{Y}]$ with uniform weights $\lambda=(1/2,1/2)$ is given by the network class $\gamma(1/2)=[(\frac{1}{2}\mathbf{X}_{b}+\frac{1}{2}\mathbf{Y}_{b},\mathrm{q}_{b})]$. Indeed, by using the triangle inequality, for every network class $\mathbb{Z}$] we have:
@@
In general, $\gamma(t)$ is a GW barycenter between $\mathbb{X}]$ and $\mathbb{Y}]$ with weights $\lambda=(1-t,t)$ because for every network class $\mathbb{Z}$] we have:
@@
Let us show that a minimizer $\mathbb{Y}^{*}]$ of a differentiable functional in GW space, $G:\mathcal{GW}\to\mathbb{R}$, must satisfy the first-order optimality condition, that is, $\nabla G([\mathbb{Y}^{*}])=0$. First, by assumption we have that $G([\mathbb{Y}^{*}])\leq G([\mathbb{Y}])$ for all $\mathbb{Y}]\in\mathcal{GW}$. In particular, $G([\mathbb{Y}^{*}])\leq G(Exp_{[\mathbb{Y}^{*}]}(tv))$ for all $v]\in\mathrm{Tan}_{[\mathbb{Y}^{*}]}$. Thus,
@@
In particular, for the direction $v]=-\nabla G([\mathbb{Y}^{*}])$ (which exists because tangent vectors are modeled as elements of $L^{2}$-spaces) we have
@@
Thus, the authors in [15] develop a first-order calculus on GW space–comprising directional derivatives of functionals, gradients as linear functionals (or vectors), first-order optimality conditions for minimizers, and linear approximations of functions around a point–by equipping it with a formal tangent space at each point $\mathbb{Y}]$, based on a Riemannian-like framework. This ends the remark.
For completeness, we revisit the proof of Proposition 4.3, originally presented in [15, Prop. 8]. Before doing so, we state the following lemma and a subsequent observation.
@@
    For any $v\in L^{2}(Y\times Y,\mu_{Y}\otimes\mu_{Y})$, $v\simeq\varphi^{*}v\in L^{2}(Z\times Z,\mu_{Z}\otimes\mu_{Z})$ (i.e., $v]=[\varphi^{*}v]\in\mathrm{Tan}_{[\mathbb{Y}]}$).